# Gaussian Splatting Lucas-Kanade

### v1

@@ A.2 Additional implementation Details
After ensuring that most Gaussians in the static regions have been sufficiently constrained, we apply a large motion mask to reconstruct small and fast-moving objects. The motion mask is defined similarly as prior work Li2021NeuralScenes as a binary segmentation mask at pixels with normalized motions larger than $0.1$. This additional motion masking step allows us to focus on rapid motions when most of the scene motions have converged.
@@ A.3 Further Discussions on Motion Learning
Geometry enhances novel view reconstruction quality. Fig. 6 presents novel view synthesis results and learned geometries from our method alongside two baseline models. This comparison highlights the impact of accurate underlying geometries on synthesized novel views. All three models converge on the training views, so we focus on the results for unseen viewpoints.
@@
Figure 6: Visualization of Gaussian trajectories and comparison of novel view synthesis results with learned geometries.
4DGaussians wu20234dgaussians learns the warp field without incorporating motion cues. This approach leads to a noisy underlying geometry with artifacts around specular areas, likely introduced to compensate for photometric inaccuracies. Consequently, its synthesized dinosaur exhibits blurriness around the edges.
@@
Another baseline predicts velocities directly instead of relying on analytical derivation as in Du et al. du2021neural . We adopt the velocity field model from Du et al. and perform time integration to derive scene flow. However, this data-driven velocity field tends to produce an "averaged" geometry, resulting in blurriness around the dinosaur’s head area.
@@
In contrast, our method generates the most crisp depth map and the highest quality synthesized result. This improvement stems from our ability to accurately learn dynamic geometries. The left-hand side of the figure demonstrates that even with monocular sequences, our method effectively separates the foreground, background, and carried object. Our method serves as a general tool for enhancing dynamic Gaussian splatting frameworks without requiring additional network modules.
@@
Motion thresholding visualizations. Fig. 7 presents a comparison of motion thresholds at different travel distances and our learned trajectories. The first row visualizes the 3D trajectories of Gaussians, color-coded by z-depth magnitude and in learned color. This visualization clearly demonstrates the accurate motions of a human waving a balloon. The second row displays the motion plot generated by our method, while the third row showcases the results from the DeformableGS baseline model Yang2023DeformableReconstruction . The baseline model exhibits a significant error, moving the stationary human instead of the balloon. Notably, the side of the balloon furthest from the human experiences the least movement, while the human’s torso exhibits the most displacement. This outcome directly contradicts expectations. In contrast, our method accurately captures the motions, correctly ranking their magnitudes. This highlights the effectiveness of our approach in distinguishing and representing the dynamic elements within the scene.
@@
Figure 7: 3D trajectories and thresholded motion plots. The first row of motion plots are results from our method and the second row are results from baseline model DeformableGS.
@@
Sensitivities to SfM initialization continued. As discussed in the limitations section, our method relies on analytical warping from canonical space, making it sensitive to Structure-from-Motion (SfM) initialization, particularly when dynamic regions lack sufficient point correspondences. Fig. 8 illustrates this sensitivity using the "train" scene from the DAVIS dataset Perazzi2016ASegmentation . The left side of the figure shows the SfM-initialized mesh, where the toy train traveling on the circular track is not captured. The top right figure displays the learned trajectories from the baseline model, exhibiting irregular Gaussian movements within the dynamic area.
@@
Figure 8: SfM initialized mesh and the reconstructed toy train trajectories. The right top subfigure shows trajectories from the baseline model, and the right bottom subfigure shows trajectories from our model.
More results on real scenes. More qualitative results on dynamic scenes can be found in Fig. 9.
Figure 9: Qualitative comparisons on Dynamic Scenes Dataset Yoon2020NovelCamera

### v2

@@ A.2 Additional implementation Details
After ensuring that most Gaussians in the static regions have been sufficiently constrained, we apply a large motion mask to reconstruct small and fast-moving objects. The motion mask is defined similarly as prior work Li et al. (2021d) as a binary segmentation mask at pixels with normalized motions larger than $0.1$. This additional motion masking step allows us to focus on rapid motions when most of the scene motions have converged.
@@ A.3 Further Discussions on Motion Learning
Geometry enhances novel view reconstruction quality. Fig. A1 presents novel view synthesis results and learned geometries from our method alongside two baseline models. This comparison highlights the impact of accurate underlying geometries on synthesized novel views. All three models converge on the training views, so we focus on the results for unseen viewpoints.
@@
Figure A1: Visualization of Gaussian trajectories and comparison of novel view synthesis results with learned geometries.
Figure A2: Qualitative comparison of results on scenes from the DyCheck dataset.
@@
4DGaussians Wu et al. (2024) learns the warp field without incorporating motion cues. This approach leads to a noisy underlying geometry with artifacts around specular areas, likely introduced to compensate for photometric inaccuracies. Consequently, its synthesized dinosaur exhibits blurriness around the edges. Another baseline predicts velocities directly instead of relying on analytical derivation as in Du et al. Du et al. (2021). We adopt the velocity field model from Du et al. and perform time integration to derive scene flow. However, this data-driven velocity field tends to produce an “averaged” geometry, resulting in blurriness around the dinosaur’s head area.
@@
In contrast, our method generates the most crisp depth map and the highest quality synthesized result. This improvement stems from our ability to accurately learn dynamic geometries. The left-hand side of the figure demonstrates that even with monocular sequences, our method effectively separates the foreground, background, and carried object. Our method serves as a general tool for enhancing dynamic Gaussian splatting frameworks without requiring additional network modules. This is further exemplified by the performance on DyCheck dataset, as shown in Fig. A2
@@
Motion thresholding visualizations. Fig. A3 presents a comparison of motion thresholds at different travel distances and our learned trajectories. The first row visualizes the 3D trajectories of Gaussians, color-coded by z-depth magnitude and in learned color. This visualization clearly demonstrates the accurate motions of a human waving a balloon. The second row displays the motion plot generated by our method, while the third row showcases the results from the DeformableGS baseline model Yang et al. (2023). The baseline model exhibits a significant error, moving the stationary human instead of the balloon. Notably, the side of the balloon furthest from the human experiences the least movement, while the human’s torso exhibits the most displacement. This outcome directly contradicts expectations. In contrast, our method accurately captures the motions, correctly ranking their magnitudes. This highlights the effectiveness of our approach in distinguishing and representing the dynamic elements within the scene.
@@
Figure A3: 3D trajectories and thresholded motion plots. The first row of motion plots are results from our method and the second row are results from baseline model DeformableGS.
@@
Sensitivities to SfM initialization continued. As discussed in the limitations section, our method relies on analytical warping from canonical space, making it sensitive to Structure-from-Motion (SfM) initialization, particularly when dynamic regions lack sufficient point correspondences. Fig. A4 illustrates this sensitivity using the “train” scene from the DAVIS dataset Perazzi et al. (2016). The left side of the figure shows the SfM-initialized mesh, where the toy train traveling on the circular track is not captured. The top right figure displays the learned trajectories from the baseline model, exhibiting irregular Gaussian movements within the dynamic area.
@@
Figure A4: SfM initialized mesh and the reconstructed toy train trajectories. The right top subfigure shows trajectories from the baseline model, and the right bottom subfigure shows trajectories from our model.
Sensitivity to the choice of depth and flow supervision Understanding the sensitivity of the choice of depth and flow supervision is crucial for accurately demonstrating the merits of our work. To explore this further, we performed experiments on the “Playground” scene in which different methods for depth and flow supervision were used. Visual comparisons of the ablated priors are shown in Fig. A5. For depth supervision, we tested Depth Anything V2 Yang et al. (2024) and Midas Birkl et al. (2023), achieving PSNR scores of 26.96 and 26.78, respectively. For flow supervision, we evaluated FlowNet Ilg et al. (2017) and MemFlow Dong & Fu (2024), achieving PSNR scores of 26.38 and 27.04. As a reference, our original setup achieved a PSNR of 26.34. The results indicate that Midas yields the poorest performance, producing blurry depth maps. In contrast, Depth Anything V2 provides sharper details but struggles with background accuracy in outdoor scenes. Our chosen method, Marigold Ke et al. (2024), gives more visually accurate depth maps, capturing even distant background details. However, it performs slightly worse quantitatively compared to the others. Nevertheless, the variation in depth priors appears to have minimal impact on our method’s final results, as most recent methods exhibit similar performance levels. Regarding flow estimation, FlowNet produced less accurate flow maps, failing to capture scene movement effectively. Conversely, MemFlow, the current state-of-the-art optical flow estimator, achieved the best results with sharp details and accurate flow. Our selected method, RAFT Teed & Deng (2020), also provides accurate flow maps but is not as sharp as MemFlow, which is likely responsible for the improved PSNR.
Figure A5: Qualitative comparison of various depth and flow methods applied to the Playground scene. Methods discussed in the paper are shown in the right column.
Network Jacobians as uncertainty measure. Deformation networks exhibit distinct patterns in regions that are easier to reconstruct compared to those that are more challenging. To explore this, we analyze the Jacobian of these deformations with respect to time, enabling us to quantify uncertainty directly from these patterns. Here in Fig. A6 we present preliminary results on a controlled synthetic scene, where the red and green balls have randomized positions in the training dataset, while the blue ball accurately corresponds to the correct time step. Brighter colors indicate higher amounts of uncertainty.
Figure A6: The uncertainties of the Gaussians are visualized in the plot, with the brighter regions higher in magnitude.
More results on real scenes. More qualitative results on dynamic scenes can be found in Fig. A7.
Figure A7: Qualitative comparisons on Dynamic Scenes Dataset Yoon et al. (2020)